OpenStudy (anonymous):

What is the exact area of a circle whose diameter is 14 cm? A. 14 straight pi cm2 B. 28 straight pi cm2 C. 49 straight pi cm2 D. 196 straight pi cm2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

c

OpenStudy (michele_laino):

here we can apply this formula: \[\huge area = \frac{{\pi {D^2}}}{4} = \frac{{\pi \times {{14}^2}}}{4} = ...?\]

OpenStudy (michele_laino):

so, I confirn your answer

OpenStudy (michele_laino):

oops..confirm*

OpenStudy (anonymous):

k

OpenStudy (michele_laino):

:)